Hi,
I'm helping my son with his homework and I myself need help understanding it.
The question is:
Find the measure of an angle if its measure is four times the measure of its complement.

Thank you for help.

Answer by ptaylor(1894)

Hi,
I'm helping my son with his homework and I myself need help understanding it.
The question is:
Find the measure of an angle if its measure is four times the measure of its complement.
Thank you for help.
Angles are said to be complementary if they add up to 90 degrees
Lex x= the angle
Then 90-x= its complement
Now we are told that the measure of the angle is 4 times its complement, so:
x=4(90-x) get rid of parens
x=360-4x add 4x to both sides
x+4x=360-4x+4x collect like terms
5x=360 divide both sides by 5
x=72 degrees ------------------answer

CK
90-72=18 degrees ------------the complement
4*18=72
72=72

Hope this helps---ptaylor